A person drops a cylindrical steel bar (Y= 1.10 x 1011) from a height of 3.50m (distance between the floor and the bottom of the vertically oriented bar). The bar, of length 0.86 m, radius 0.55 cm, and mass 1.70 kg, hits the floor and bounces up, maintaining its vertical orientation. Assuming the collision with the floor is elastic, and that no rotation occurs, what is the maximum compression of the bar? (answer in mm)
Answer #1

You can the velocity of the bar when it hits the ground using kinematics:

v² = vo² + 2 a d

v² = 2 a d

Calculate kinetic energy:

KE = 1/2 m v²

KE = 1/2 m (2 a d)

KE = m a d

Potential energy:

PE = 1/2 k x²

PE = 1/2 Y A x² / L

PE = 1/2 Y A x² / L

PE = 1/2 Y π R² x² / L

Set potential energy equal to kinetic energy and solve for x:

PE = KE

1/2 Y π R² x² / L = m a d

x = √{2 m a d L / [Y π R²]}

x = √{2 (1.70 kg) (-9.81 m/s²) (-3.50 m) (0.86 m) / [(1.10 * 10^11 Pa) π (0.0055 m)²]}

x = 3.10 * 10^-3 m

x = 3.10 mm